# Low evidence for ringdown overtone in GW150914 when marginalizing over time and sky location uncertainty

### v3

@@ Abstract
Tests of the no-hair theorem using astrophysical black holes involve the detection of at least two quasi-normal modes (QNMs) in the gravitational wave emitted by a perturbed black hole. A detection of two modes — the dominant, $(\ell,m,n)=(2,2,0)$, mode and its first overtone, the $(2,2,1)$ mode — in the post-merger signal of the binary black hole merger GW150914 was claimed in Isi et al. [1], with further evidence provided in Isi & Farr [2]. However, Cotesta et al. [3] disputed this claim, finding that evidence for the overtone only appeared if the signal was analyzed before merger, when a QNM description of the signal is not valid. Due to technical challenges, both of these analyses fixed the merger time and sky location of GW150914 when estimating the evidence for the overtone. At least some of the contention can be attributed to fixing these parameters. Here, we surmount this difficulty and fully marginalize over merger time and sky-location uncertainty while doing a post-merger QNM analysis of GW150914. We find that marginalizing over all parameters yields low evidence for the presence of the overtone, with a Bayes factor of $1.10\pm 0.03$ in favor of a QNM model with the overtone versus one without. The arrival time uncertainty of GW150914 is too large to definitively claim detection of the $(2,2,1)$ mode.
@@ I Introduction
The direct detection of gravitational waves has made it possible to perform “Black hole spectroscopy” on real astrophysical black holes. Black holes that are formed from a binary black hole (BBH) merger are initially perturbed. The perturbation is radiated away in a gravitational-wave “ringdown” as the black hole settles down to a Kerr space time. The ringdown is a superposition of damped sinusoids, or quasi-normal modes (QNMs), that are uniquely determined by three indices: two angular indices $\ell\geq 2$ and $m\leq|\ell|$, and an overtone number $n\geq 0$ [4, 5]. As a consequence of the no-hair theorem, we expect that the frequencies and damping times of all the QNMs are fully characterized by the final black hole’s mass and spin. The Kerr nature of the final object can be tested if more than one QNM can be observed in a post-merger gravitational wave (GW) [6]. Should the QNMs be found to need more than two parameters to characterize their frequency and damping times it would indicate that the final object is not a Kerr black hole as expected from General Relativity (GR).
@@
Black hole spectroscopy is challenging to accomplish in practice. The QNMs are exponentially damped, and so the signal-to-noise ratio (SNR) of the signal falls off quickly after merger. This would favor searching for QNMs immediately after merger, but it was thought that one had to wait $\sim 10\,M$ (where $M$ is the mass of the black hole, in seconds) or longer (depending on the sensitivity of the detectors [7, 8]) after the time at which the GW amplitude peaks ($t_{c}$) for non-linearities to radiate away and a QNM description of the signal to be valid [9, 10, 11, 12, 13]. The $10\,M$ rule was arrived at by studying the stability of the “fundamental” $n=0$ modes in numerical signals. However, in Giesler et al. [14] it was shown that including overtones of the dominant, $(\ell,m,n)=(2,2,0)$ mode — i.e., modes with $(2,2,n\geq 1)$ — yielded surprisingly good fits to a numerical relativity (NR) simulation at $t_{c}$. This led Giesler et al. to claim that the spacetime could be described by a linearly perturbed black hole at the merger. Isi et al. [1] then performed an overtone analysis of the first detected binary black hole merger, GW150914 [15], and claimed detection of the $(2,2,1)$ mode at $3.6\,\sigma$. They performed a no-hair test using the overtone by allowing its frequency and damping time to vary in a Bayesian analysis, finding that it was broadly consistent with GR.
@@
The validity of using a QNM description of the signal at $t_{c}$ is controversial. Although it was known that overtones yielded a better fit to the GW at merger prior to Giesler et al. [9], it was not clear whether this indicated that the spacetime was truly described by a linearly perturbed BH (as claimed in Giesler et al.) or if it was simply due to the overtones overfitting the signal (meaning that any sufficiently compatible signal could be fit by them). While Giesler et al. provided some evidence for the former, subsequent analyses have challenged their interpretation [16, 17, 18, 19, 20, 21, 22], leaving the question unresolved.
@@
Even the claim of a detection of an overtone in GW150914 by Isi et al. [1] has been challenged. Cotesta et al. [3] performed an overtone analysis at multiple times around the purported $t_{c}$ of GW150914. At each time they calculated the Bayes factor for a signal model consisting of the $(2,2,0){+}(2,2,1)$ QNMs versus one consisting of just the $(2,2,0)$ mode. Assuming equal prior weight for the two models, the Bayes factor $\mathcal{B}=Z_{220{+}221}/Z_{220}$ (where $Z_{x}$ is the Bayesian evidence for model $x$) gives the odds ratio that the overtone model is a better description of the signal. Cotesta et al. found that the Bayes factor was $\sim 1$ at $t_{c}$, indicating that the data were uninformative as to the presence of the $(2,2,1)$ mode. They claimed that positive evidence for the overtone only appeared if the QNM model was used prior to the merger. Since a QNM model is not valid there, this would indicate that the overtone model was simply overfitting the data. The low Bayes factor was also consistent with what had been reported in Bustillo et al. [23] and in publications from the LIGO and Virgo Collaborations (LVC) [24]. [^1] It should be noted that the pipeline  [38, 39] used to report overtone Bayes factors by the LVC is the same as used in Cotesta et al., so the agreement between the two is not entirely surprising. However, in a rebuttal, Isi & Farr [2] found that the Bayes factor for the overtone was nearly $100$ at $t_{c}$, reinforcing the initial finding in Isi et al. [1].
@@
The disagreement between Isi et al. and Cotesta et al. may be due to several factors, including data conditioning choices and the manner in which the likelihood function is evaluated. The effects of these choices are investigated in Wang et al. [26]. It is evident, however, that the choice of start time for the ringdown is one of the most important factors — if not the defining factor [27] — in determining if the overtone is present. While they may disagree on the magnitude of the Bayes factor at a given time, both Cotesta et al. and Isi & Farr found that its value changes rapidly as the ringdown start time is varied around $t_{c}$, increasing by several orders of magnitude in less than $5M$. For GW150914, $5M$ corresponds to $\sim 1.7\,$ms [15]. This is strikingly small; for comparison, the statistical uncertainty on $t_{c}$ when using an inspiral-merger-ringdown (IMR) template to model the full signal is $\sim 1.2\,$ms [28, 29]
@@
The statistical uncertainty in $t_{c}$ arises from an overall uncertainty in the arrival time of the signal in a single detector, and the uncertainty in the sky location of GW150914. The latter affects the relative offset in arrival times between the Hanford and Livingston detectors, which observed the event [15]. A full accounting of the evidence for or against the overtone should marginalize over the time and sky location uncertainty of GW150914. However, the analysis pipelines used in Isi et al. and Cotesta et al. fix the arrival time and sky location of the signal to their maximum likelihood values from IMR studies.
@@
Finch & Moore were able to marginalize over the sky location and arrival time in their post-merger analysis of GW150914 [35]. They accomplished this by attaching an arbitrary signal model to the start of the ringdown. This had the effect of smoothly tapering the signal model to zero, thereby allowing them to use the traditional frequency-domain likelihood function that is used in standard parameter estimation analyses. Doing so, they recovered Bayes factors $B_{220}^{221}$ between 0.1 and 10, depending on the prior chosen on $t_{c}$. Their prior choices consisted of Gaussian distributions of width $1M$ centered at various $t_{c}$ values.
@@
By using the traditional frequency-domain likelihood function Finch & Moore coupled information from the pre-merger part of the signal with the post-merger QNM model. The frequency-domain likelihood is equivalent to the convolution of the “whitened” data with the “whitened” signal model. The whitening filter couples adjacent points in time over several milliseconds. This means, for example, that the beginning of the whitened ringdown will be a weighted sum over a time span starting before the merger that is on order with the damping time of the overtone ($\sim 1.4\,$ms). Consequently, the recovered QNM parameters will be influenced by both the choice of signal model and the data just prior to the start of the ringdown.
@@
Here, we overcome the technical hurdles that arise when marginalizing over time and sky location uncertainty while doing a purely QNM analysis of the post-merger portion of GW150914. We simultaneously analyze the pre-merger data using the inspiral-merger portion of an IMR template waveform. The pre- and post-merger portions of the analysis only share a common sky location and geocentric $t_{c}$. This allows us to use the full signal to isolate the binary, giving the most accurate assessment of the timing uncertainty, while still maintaining independence between the pre- and post-merger analyses. In a companion paper [36] we show that it is necessary to model the entire observable signal in order to marginalize over timing uncertainties. Consequently, our analysis here is the most complete accounting of all uncertainties that can be preformed when using a QNM-only model of the post-merger signal.
@@
Figure 1: Whitened strain in the Hanford (top) and Livingston (bottom) detectors around GW150914. Plotted is the original data (black line) along with the maximum likelihood waveform, which consists of an inspiral portion (blue lines) and ringdown portion (red lines). The inspiral and ringdown models are independent of each other, and only share a common sky location and geocentric $t_{c}$. Ringdown (inspiral) waveforms are gated and in-painted before (after) the $t_{c}$ in each detector; the first (last) non-zero point in the waveform is indicated by the solid dot. The data are also gated and in-painted before (after) the detector $t_{c}$ for the ringdown (inspiral) model, so that neither model simultaneously includes the same data point from the signal in the likelihood. Individual waveforms drawn from the upper $90$th percentile of the posterior are also plotted to indicate the uncertainty region. Vertical purple lines indicate the gate time for each of the waveforms plotted. The time axis is zeroed with respect to the maximum likelihood $t_{c}$ in each detector $t_{\rm ref}^{H/L}$ from the 4-OGC IMR analysis [29].
@@ II Methods
An in-depth description of our analysis is presented in a companion paper [36]. Here, we give a brief overview.
@@
There are two approaches around this. One is to numerically invert $\bm{\Sigma}_{D}$ and explicitly calculate the matrix multiplication in Eq. (1). This is what is done by the “ringdown” [37] and “pyring” [38, 39] analyses used by Isi et al. and Cotesta et al., respectively. Numerically inverting $\bm{\Sigma}_{D}$ is computationally expensive, but need only to be done once per analysis if the start time of the ringdown does not change. If the start time of the ringdown does change (meaning different times need to be excised from the data), then $\bm{\Sigma}^{-1}_{D}$ and $\det\bm{\Sigma}_{D}$ need to be recalculated, making the analysis substantially more expensive.
The other approach is to use “gating and in-painting” to solve the likelihood [40]. This is mathematically equivalent to the brute force numerical method, but is more efficient for longer stretches of data [41, 26]. In this approach, a “gate” is applied to a stretch of data, within which the data is zeroed out. The gated portion is then “in-painted” so that it contributes nothing to the likelihood. For a ringdown-only analysis the gate ends at the start of the ringdown; we use a 111\,1s gate to ensure that the rest of the observable signal is removed. The gate shifts around in time as the start time of the ringdown is varied, thereby removing more or less of the signal from the analysis.
@@
We also show in Ref. [36] that it is not possible to only model part of a signal when marginalizing over $t_{c}$ and sky location. Marginalizing over these parameters while only modelling part of the signal means that the portion of data that is excised from the analysis is allowed to shift around in time. If the shift is such that the entire observable signal is excised from the analysis, then a relatively large likelihood can be obtained for any parameter value by reducing the overall amplitude of the template to noise level. Consequently, the posterior probability will favor excising as much of the signal as possible, even the part that the template is meant to model. This means that for a ringdown-only analysis the start time of the template will be pushed to as late as the prior will allow. Likewise, if an inspiral-only analysis is performed (in which the post-merger is excised and the pre-merger is modelled with a template) then the end time of the template will be pushed to as early as the prior will allow. It is necessary to have a signal model for all times that a signal is observable in the data when doing any sort of marginalization over arrival time.
@@
We overcome this problem by simultaneously yet independently modelling the pre- and post-merger data. We treat the pre- and post-merger signal as separate analyses that share a common set of hyperparameters — the sky location and geocentric $t_{c}$. We use gating and in-painting for both. For the pre-merger analysis, we excise the data after $t_{c}$ while for the post-merger analysis we excise the data before $t_{c}$. The pre-merger portion of the signal is modelled using the IMRPhenomXPHM approximant [43], which we zero out for all times after $t_{c}$. The post-merger signal is modelled using a QNM model consisting of the $(2,2,0){+}(2,2,1)$ modes, which is zeroed out for all times before $t_{c}$. The final mass and final spin in the QNM model are varied over uniform priors such that $M_{f}\in[10,200]M_{\odot}$ and $\chi_{f}\in[-0.99,0.99]$. For the coalescence time we use a uniform prior spanning $t_{\rm ref}\pm 0.1\,$s, where $t_{\rm ref}$ is the maximum likelihood geocentric coalescence time from Ref. [29]. This is the same prior width on $t_{c}$ as used in standard IMR analyses.
@@
For the $(2,2,0)$ amplitude $A_{220}$ we sample over the range $[10^{-25},8\times 10^{-17}]$ uniformly in log base 10. For the $(2,2,1)$ amplitude $A_{221}$ we use a uniform prior on $A_{221}/A_{220}\in[0,5)$. We additionally impose the constraint that the $(2,2,0)$ mode’s contribution to the post-merger SNR be $>2$. This is to ensure that at least the fundamental mode is present in the post-merger signal. Otherwise, we obtain samples where the $(2,2,1)$ mode appears to have large amplitude, but is in fact matching the $(2,2,0)$ mode in the signal (the “label switching” problem).
@@
Gating the pre- and post-merger means that no coupling occurs over the boundary due to whitening. Simultaneously modelling both portions has the added benefit that we use the entire signal to constrain the sky location. We use PyCBC Inference [44] with the dynesty sampler [45] to perform the Bayesian analysis.

### v4

@@ Abstract
Tests of the no-hair theorem using astrophysical black holes involve the detection of at least two quasinormal modes (QNMs) in the gravitational wave emitted by a perturbed black hole. A detection of two modes—the dominant, $(\ell,m,n)=(2,2,0)$, mode and its first overtone, the $(2,2,1)$ mode—in the postmerger signal of the binary black hole merger GW150914 was claimed by Isi et al. [1], with further evidence provided by Isi & Farr [2]. However, Cotesta et al. [3] disputed this claim, finding that evidence for the overtone only appeared if the signal was analyzed before merger, when a QNM description of the signal is not valid. Because of technical challenges, both of these analyses fixed the merger time and sky location of GW150914 when estimating the evidence for the overtone. At least some of the contention can be attributed to fixing these parameters. Here, we surmount this difficulty and fully marginalize over merger time and sky-location uncertainty while doing a postmerger QNM analysis of GW150914. We find that marginalizing over all parameters yields low evidence for the presence of the overtone, with a Bayes factor of $1.10\pm 0.03$ in favor of a QNM model with the overtone versus one without. The arrival time uncertainty of GW150914 is too large to definitively claim detection of the $(2,2,1)$ mode.
@@ I Introduction
The direct detection of gravitational waves has made it possible to perform “black hole spectroscopy” on real astrophysical black holes. Black holes that are formed from a binary black hole (BBH) merger are initially perturbed. The perturbation is radiated away in a gravitational-wave “ringdown” as the black hole settles down to a Kerr spacetime. The ringdown is a superposition of damped sinusoids, or quasinormal modes (QNMs), that are uniquely determined by three indices: two angular indices $\ell\geq 2$ and $m\leq|\ell|$, and an overtone number $n\geq 0$ [4, 5]. As a consequence of the no-hair theorem, we expect that the frequencies and damping times of all the QNMs are fully characterized by the final black hole’s mass and spin. The Kerr nature of the final object can be tested if more than one QNM can be observed in a postmerger gravitational wave (GW) [6]. Should the QNMs be found to need more than two parameters to characterize their frequency and damping times it would indicate that the final object is not a Kerr black hole as expected from General Relativity (GR).
@@
Black hole spectroscopy is challenging to accomplish in practice. The QNMs are exponentially damped, and so the signal-to-noise ratio (SNR) of the signal falls off quickly after merger. This would favor searching for QNMs immediately after merger, but it was thought that one had to wait $\sim 10M$ (where $M$ is the mass of the black hole, in seconds) or longer (depending on the sensitivity of the detectors [7, 8]) after the time at which the GW amplitude peaks ($t_{c}$) for nonlinearities to radiate away and a QNM description of the signal to be valid [9, 10, 11, 12, 13]. The $10M$ rule was arrived at by studying the stability of the “fundamental” $n=0$ modes in numerical signals. However, Giesler et al. [14] show that including overtones of the dominant, $(\ell,m,n)=(2,2,0)$ mode—i.e., modes with $(2,2,n\geq 1)$—yielded surprisingly good fits to a numerical relativity simulation at $t_{c}$. This led Giesler et al. to claim that the spacetime could be described by a linearly perturbed black hole at the merger. Isi et al. [1] then performed an overtone analysis of the first detected binary black hole merger, GW150914 [15], and claimed detection of the $(2,2,1)$ mode at $3.6\,\sigma$. They performed a no-hair test using the overtone by allowing its frequency and damping time to vary in a Bayesian analysis, finding that it was broadly consistent with GR.
@@
The validity of using a QNM description of the signal at $t_{c}$ is controversial. Although it was known prior to the work of Giesler et al. that overtones yielded a better fit to the GW at merger  [9], it was not clear whether this indicated that the spacetime was truly described by a linearly perturbed black hole (as claimed in Giesler et al.) or if it was simply due to the overtones overfitting the signal (meaning that any sufficiently compatible signal could be fit by them). While Giesler et al. provided some evidence for the former, subsequent analyses have challenged their interpretation [16, 17, 18, 19, 20, 21, 22], leaving the question unresolved.
@@
Even the claim of a detection of an overtone in GW150914 by Isi et al. [1] has been challenged. Cotesta et al. [3] performed an overtone analysis at multiple times around the purported $t_{c}$ of GW150914. At each time they calculated the Bayes factor for a signal model consisting of the $(2,2,0){+}(2,2,1)$ QNMs versus one consisting of just the $(2,2,0)$ mode. Assuming equal prior weight for the two models, the Bayes factor $\mathcal{B}=Z_{220{+}221}/Z_{220}$ (where $Z_{x}$ is the Bayesian evidence for model $x$) gives the odds ratio that the overtone model is a better description of the signal. Cotesta et al. found that the Bayes factor was $\sim 1$ at $t_{c}$, indicating that the data were uninformative as to the presence of the $(2,2,1)$ mode. They claimed that positive evidence for the overtone only appeared if the QNM model was used prior to the merger. Since a QNM model is not valid there, this would indicate that the overtone model was simply overfitting the data. The low Bayes factor was also consistent with what had been reported by Bustillo et al. [23] and in publications from the LIGO and Virgo Collaborations (LVC) [24] [^1] It should be noted that the pipeline [38, 39] used to report overtone Bayes factors by the LVC is the same as used in Cotesta et al., so the agreement between the two is not entirely surprising.. However, in a rebuttal, Isi and Farr [2] found that the Bayes factor for the overtone was nearly $100$ at $t_{c}$, reinforcing the initial finding by Isi et al. [1].
@@
The disagreement between Isi et al. and Cotesta et al. may be due to several factors, including data conditioning choices and the manner in which the likelihood function is evaluated. The effects of these choices were investigated by Wang et al. [26]. It is evident, however, that the choice of start time for the ringdown is one of the most important factors—if not the defining factor [27]—in determining if the overtone is present. While they may disagree on the magnitude of the Bayes factor at a given time, both Cotesta et al. and Isi and Farr found that its value changes rapidly as the ringdown start time is varied around $t_{c}$, increasing by several orders of magnitude in less than $5M$. For GW150914, $5M$ corresponds to $\sim 1.7\,$ms [15]. This is strikingly small; for comparison, the statistical uncertainty on $t_{c}$ when using an inspiral-merger-ringdown (IMR) template to model the full signal is $\sim 1.2\,$ms [28, 29]
@@
The statistical uncertainty in $t_{c}$ arises from an overall uncertainty in the arrival time of the signal in a single detector, and the uncertainty in the sky location of GW150914. The latter affects the relative offset in arrival times between the Hanford and Livingston detectors, which observed the event [15]. A full accounting of the evidence for or against the overtone should marginalize over the time and sky location uncertainty of GW150914. However, the analysis pipelines used by Isi et al. and Cotesta et al. fix the arrival time and sky location of the signal to their maximum likelihood values from IMR studies.
@@
Finch and Moore were able to marginalize over the sky location and arrival time in their postmerger analysis of GW150914 [35]. They accomplished this by attaching an arbitrary signal model to the start of the ringdown. This had the effect of smoothly tapering the signal model to zero, thereby allowing them to use the traditional frequency-domain likelihood function that is used in standard parameter estimation analyses. Doing so, they recovered Bayes factors $B_{220}^{221}$ between 0.1 and 10, depending on the prior chosen on $t_{c}$. Their prior choices consisted of Gaussian distributions of width $1M$ centered at various $t_{c}$ values.
@@
By using the traditional frequency-domain likelihood function, Finch and Moore coupled information from the premerger part of the signal with the postmerger QNM model. The frequency-domain likelihood is equivalent to the convolution of the “whitened” data with the whitened signal model. The whitening filter couples adjacent points in time over several milliseconds. This means, for example, that the beginning of the whitened ringdown will be a weighted sum over a time span starting before the merger that is on order with the damping time of the overtone ($\sim 1.4\,$ms). Consequently, the recovered QNM parameters will be influenced by both the choice of signal model and the data just prior to the start of the ringdown.
@@
Here, we overcome the technical hurdles that arise when marginalizing over time and sky location uncertainty while doing a purely QNM analysis of the postmerger portion of GW150914. We simultaneously analyze the premerger data using the inspiral-merger portion of an IMR template waveform. The pre and postmerger portions of the analysis only share a common sky location and geocentric $t_{c}$. This allows us to use the full signal to isolate the binary, giving the most accurate assessment of the timing uncertainty, while still maintaining independence between the pre and postmerger analyses. In a companion paper [36] we show that it is necessary to model the entire observable signal in order to marginalize over timing uncertainties. Consequently, our analysis here is the most complete accounting of all uncertainties that can be preformed when using a QNM-only model of the postmerger signal.
@@
Figure 1: Whitened strain in the Hanford (top) and Livingston (bottom) detectors around GW150914. Plotted is the original data (black line) along with the maximum likelihood waveform, which consists of an inspiral portion (blue lines) and ringdown portion (red lines). The inspiral and ringdown models are independent of each other, and only share a common sky location and geocentric $t_{c}$. Ringdown (inspiral) waveforms are gated and in-painted before (after) the $t_{c}$ in each detector; the first (last) nonzero point in the waveform is indicated by the solid dot. The data are also gated and in-painted before (after) the detector $t_{c}$ for the ringdown (inspiral) model, so that neither model simultaneously includes the same data point from the signal in the likelihood. Individual waveforms drawn from the upper $90$th percentile of the posterior are also plotted to indicate the uncertainty region. Vertical purple lines indicate the gate time for each of the waveforms plotted. The time axis is zeroed with respect to the maximum likelihood $t_{c}$ in each detector $t_{\rm ref}^{H/L}$ from the 4-OGC IMR analysis [29].
@@ II Methods
An in-depth description of our analysis is presented in a following paper [36]. Here, we give a brief overview.
@@
There are two approaches around this. One is to numerically invert $\bm{\Sigma}_{D}$ and explicitly calculate the matrix multiplication in Eq. (1). This is what is done by the “ringdown” [37] and “pyring” [38, 39] analyses used by Isi et al. and Cotesta et al., respectively. Numerically inverting $\bm{\Sigma}_{D}$ is computationally expensive but needs only to be done once per analysis if the start time of the ringdown does not change. If the start time of the ringdown does change (meaning different times need to be excised from the data), then $\bm{\Sigma}^{-1}_{D}$ and $\det\bm{\Sigma}_{D}$ need to be recalculated, making the analysis substantially more expensive.
The other approach is to use “gating and in-painting” to solve the likelihood [40]. This is mathematically equivalent to the brute force numerical method, but is more efficient for longer stretches of data [41, 26]. In this approach, a gate is applied to a stretch of data, within which the data are zeroed out. The gated portion is then in-painted so that it contributes nothing to the likelihood. For a ringdown-only analysis the gate ends at the start of the ringdown; we use a 111\,1s gate to ensure that the rest of the observable signal is removed. The gate shifts around in time as the start time of the ringdown is varied, thereby removing more or less of the signal from the analysis.
@@
We also show in Ref. [36] that it is not possible to only model part of a signal when marginalizing over $t_{c}$ and sky location. Marginalizing over these parameters while only modeling part of the signal means that the portion of data that is excised from the analysis is allowed to shift around in time. If the shift is such that the entire observable signal is excised from the analysis, then a relatively large likelihood can be obtained for any parameter value by reducing the overall amplitude of the template to noise level. Consequently, the posterior probability will favor excising as much of the signal as possible, even the part that the template is meant to model. This means that for a ringdown-only analysis the start time of the template will be pushed to as late as the prior will allow. Likewise, if an inspiral-only analysis is performed (in which the postmerger is excised and the premerger is modeled with a template) then the end time of the template will be pushed to as early as the prior will allow. It is necessary to have a signal model for all times that a signal is observable in the data when doing any sort of marginalization over arrival time.
@@
We overcome this problem by simultaneously yet independently modeling the pre and postmerger data. We treat the pre and postmerger signal as separate analyses that share a common set of hyperparameters—the sky location and geocentric $t_{c}$. We use gating and in-painting for both. For the premerger analysis, we excise the data after $t_{c}$ while for the postmerger analysis, we excise the data before $t_{c}$. The premerger portion of the signal is modeled using the IMRPhenomXPHM approximant [43], which we zero out for all times after $t_{c}$. The postmerger signal is modeled using a QNM model consisting of the $(2,2,0){+}(2,2,1)$ modes, which is zeroed out for all times before $t_{c}$. The final mass and final spin in the QNM model are varied over uniform priors such that $M_{f}\in[10,200]M_{\odot}$ and $\chi_{f}\in[-0.99,0.99]$. For the coalescence time we use a uniform prior spanning $t_{\rm ref}\pm 0.1\,$s, where $t_{\rm ref}$ is the maximum likelihood geocentric coalescence time from Ref. [29]. This is the same prior width on $t_{c}$ as used in standard IMR analyses.
@@
For the $(2,2,0)$ amplitude $A_{220}$ we sample over the range $[10^{-25},8\times 10^{-17}]$ uniformly in log base 10. For the $(2,2,1)$ amplitude $A_{221}$ we use a uniform prior on $A_{221}/A_{220}\in[0,5)$. We additionally impose the constraint that the $(2,2,0)$ mode’s contribution to the postmerger SNR be $>2$. This is to ensure that at least the fundamental mode is present in the postmerger signal. Otherwise, we obtain samples where the $(2,2,1)$ mode appears to have large amplitude, but is in fact matching the $(2,2,0)$ mode in the signal (the “label switching” problem).
@@
Gating the pre and postmerger means that no coupling occurs over the boundary due to whitening. Simultaneously modeling both portions has the added benefit that we use the entire signal to constrain the sky location. We use PyCBC Inference [44] with the dynesty sampler [45] to perform the Bayesian analysis.